OpenStudy (anonymous):

Please please please help my time is almost up for the test.... Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10

OpenStudy (anonymous):

I'd probably use substitution because you can easily seperate an x and a y. I don't know if that is correct though. Are you sure you can use the internet for tests lol?

OpenStudy (anonymous):

Yess, we can. I asked to make sure. I dont have a clue how to answer any of this or where to start.

OpenStudy (anonymous):

Substitution works like this. You get an x buy itself, eg x = (blah) then in another problem you substitute in the (blah) instead of the x. So, x = 5y - 2z Then the second one can be (5y - 2z) + 4y - z = 12 call this (4)

OpenStudy (anonymous):

Then you need another sum so change the 2nd on and substititute it into the thrid one

OpenStudy (anonymous):

Ohhh, I get it..... i think

OpenStudy (anonymous):

x = 12 - 4y + z put it int n the next one 2(12- 4y + z) - y + 3z = 10 call this (5) now use (4) amd (5) you can solve by eliminating or substituting again.

OpenStudy (anonymous):

See now you have a sum you got rid of the x 2(12 - 4y + z) - y + 3z = 10 (5y - 2z) + 4y - z = 12 Do the same thing to get rid of another letter

OpenStudy (anonymous):

Ohh i get it.

OpenStudy (anonymous):

24 - 8y + z - y + 3z = 10 goes to 24-9y+4z = 10 and 9y-3z = 12 Oh good I'm glad it takes many steps lol

OpenStudy (anonymous):

Now add the 2 sums together because here it is too ahrd to substitute. I change the (4) to -9y+4z = 10-24 z = -2

OpenStudy (anonymous):

Thank you so much, I finally understand it

OpenStudy (anonymous):

No problem :)